OpenStudy (anonymous):

Please help? :) Aunt Bea has 37 pigs and ducks on her farm. In all, there are 116 legs. How many pigs and how many ducks does Aunt Bea have on her farm?

OpenStudy (lucaz):

each pig and duck has 2 legs 37(2)+ducks(2)=116 2d=116-74 d=42/2=21 ducks